OpenStudy (anonymous):

find the derivative f(x)= ln (x^2 square(4x+1))/(x^3+5)^3

zepdrix (zepdrix):

Hey :) Oh my, this is quite an ugly problem :o

zepdrix (zepdrix):

Log is just of the top part?

zepdrix (zepdrix):

\[\Large\sf \frac{\ln(x^2 \sqrt{4x+1})}{(x^3+5)^3}\]Was the square(4x+1) supposed to be square root as I've written it?

OpenStudy (anonymous):

\[\ln \frac{ x ^{3\sqrt{4x+1}} }{ ^{} }\]

OpenStudy (anonymous):

under (x^3+5)^3